# STATE ToxiCN: A Benchmark for Span-level Target-Aware Toxicity Extraction in Chinese Hate Speech Detection

### v2

@@ 1 Introduction
Moreover, we systematically summarize and annotate commonly used hateful slang from real-world online forums to address the challenges of identifying hateful slang. We compile a comprehensive annotated lexicon, labeling each hateful slang with its frequent groups and providing detailed explanations of its usage and contextual meaning. This is the first annotated lexicon of Chinese hateful slang with interpretative annotations. This resource not only helps to understand how hateful slang are used to disguise hate speech but also provides valuable annotated data for evaluating and improving the ability of LLMs to detect Chinese hateful slang. The main contributions of this work are summarized as follows:
@@ 2 Related Work
Hate speech detection is a critical task in NLP that has attracted considerable attention recently. Researchers have increasingly turned to pre-trained models to address this issue Caselli et al. (2020); Hanu and Unitary team (2020); Zhou et al. (2021); AlKhamissi et al. (2022a); Ali et al. (2022). To facilitate progress in this field, several datasets tailored to hate speech detection have been developed Waseem and Hovy (2016); Davidson et al. (2017); Founta et al. (2018); Hartvigsen et al. (2022). Pavlopoulos et al. (2021) introduced span-level hate speech detection, while the TBO dataset advanced the field by pioneering the extraction of Target-Argument-Harmful triples Zampieri et al. (2023). However, Chinese hate speech detection remains significantly underdeveloped.
@@ 3.3.1 Annotation Guidelines
To mitigate bias, we assemble annotators with diverse backgrounds, including differences in gender, age, ethnicity, region, and educational level. We believe this diversity helps reduce bias from a single perspective and enhances the reliability of the annotation results. All annotators possess linguistic expertise and have undergone systematic training. During the annotation process, we primarily employ regular cross-validation and expert arbitration to ensure the objectivity and consistency of annotation results. Additionally, we maintain an online document to record and update Chinese hateful slang identified in the posts. This annotated lexicon serves not only as a research resource but also helps to align annotators’ standards.
@@ 4.1 Baselines
Open-source Models: mT5-base Xue (2020), Mistral-7B Jiang et al. (2023), LLaMA3-8B AI@Meta (2024), Qwen2.5-7B Team (2024); Safety-domain Models: ShieldLM-14B-Qwen Zhang et al. (2024), ShieldGemma-9B Zeng et al. (2024), and Closed-source LLMs: LLaMA3-70B AI@Meta (2024), Qwen2.5-72B Team (2024), Gemini-1.5-Pro Team et al. (2023), Claude-3.5-Sonnet Anthropic (2024), GPT-4o OpenAI (2024), DeepSeek-v3 Liu et al. (2024).
@@ 5.3.1 Impact of Hateful Slang on Chinese Hate Speech Detection
LLM APIs demonstrate superior performance on the subset of the test set containing hateful slang. With the exception of a slight decrease in the soft matching metric for argument identification, nearly all other metrics show an improvement. We believe that this is primarily due to two factors: First, hateful slang often appears as the target, and the explicit target terms effectively compensate for the limitations of LLM APIs in span boundary recognition. Second, the identification of hateful slang often requires rich background knowledge; for example, Example 2 in the figure requires knowledge of Chinese history and related common usages, which is precisely where LLM APIs excel. Therefore, hateful slang, which is difficult for fine-tuned models to leverage, becomes an effective source of information for LLM APIs.
@@ 6 Conclusions
Our Annotated Lexicon covers a wide range of commonly used Chinese hateful slang; however, due to the dynamic nature and rapid evolution of the Chinese internet language, certain emerging or less widely recognized cloaking expressions (e.g., more complex word transformations or domain-specific slang) might not be fully captured. Finally, this study focuses primarily on textual features and does not consider non-textual elements, such as images, videos, or metadata about the authors, which may limit the model’s ability to comprehensively detect multimodal hate speech.
In the future, we aim to expand the scope of our dataset to include more diverse contexts and cloaking language, while exploring multimodal and cross-domain hate speech detection methods to improve overall performance and applicability.
@@ Appendix A Details of the Annotated Lexicon
We use JSON files to store the annotated lexicon, which is a mainstream coding specification to facilitate machine-readability. The data structure is defined as Sample = (Term, [Groups], Definition), where Groups indicates the groups that are commonly targeted by the hateful slang. If a term does not have a specific targeted group, then Groups is set to "others". Here we provide four samples in Figure 2.
Figure 2: Examples of annotated Chinese hateful slang, including Term, Groups and Definition.
@@ Appendix E Derivative Rules of Chinese Hateful Slang
In this section, we further explain hateful slang in the lexicon and analyze their hate semantics and derivative logic through classification. Lu et al. (2023) propose seven types of derivative rules: Deformation (separating and combining individual characters), Homophonic (similar pronunciation), Irony (positive words used ironically to insult), Abbreviation (shortening and contracting sensitive words), Metaphor (degrading targets into something sarcastic), Code Mixing (use of non-Chinese codes to emphasize tone), and Borrowed Words (foreign phonetic words with specific toxic cultural connotations). We add two more derivative rules, Historical Allusions and Stereotypes.

### v3

@@ 1 Introduction
Moreover, we summarize and annotate commonly used hateful slang from real-world online forums to address the challenges of identifying hateful slang. We compile a comprehensive annotated lexicon, labeling each hateful slang with its frequent groups and providing explanations of its usage and contextual meaning. This is the first annotated lexicon of Chinese hateful slang with interpretative annotations. This resource not only helps to understand how hateful slang is used to disguise hate speech, but also provides valuable annotated data to evaluate and improve the ability of LLMs to detect hateful slang. The main contributions of this work are summarized as follows:
@@ 2 Related Work
Hate speech detection is a critical task in Natural Language Processing (NLP) that has attracted considerable attention recently. Researchers have increasingly turned to pre-trained models to address this issue Caselli et al. (2020); Hanu and Unitary team (2020); Zhou et al. (2021); AlKhamissi et al. (2022a); Ali et al. (2022). To facilitate progress in this field, several datasets tailored to hate speech detection have been developed Waseem and Hovy (2016); Davidson et al. (2017); Founta et al. (2018); Hartvigsen et al. (2022). Pavlopoulos et al. (2021) introduced span-level hate speech detection, while the TBO dataset advanced the field by pioneering the extraction of Target-Argument-Harmful triples Zampieri et al. (2023). However, Chinese hate speech detection remains significantly underdeveloped.
@@ 3.3.1 Annotation Guidelines
To mitigate bias, we assemble annotators with diverse backgrounds, including differences in gender, age, ethnicity, region, and educational level. The detailed information of annotators can be found in Appendix G. All annotators possess linguistic expertise and have undergone systematic training. During the annotation process, we primarily employ regular cross-validation and expert arbitration to ensure the objectivity and consistency of annotation results. Additionally, we maintain an online document to record and update Chinese hateful slang identified in the posts. This annotated lexicon serves not only as a research resource but also helps to align annotators’ standards.
@@ 4.1 Baselines
Open-source Models: mT5-base Xue (2020), Mistral-7B Jiang et al. (2023), LLaMA3-8B AI@Meta (2024), Qwen2.5-7B Team (2024); Safety-domain Models: ShieldLM-14B-Qwen Zhang et al. (2024), ShieldGemma-9B Zeng et al. (2024), and Closed-source LLMs: LLaMA3-70B AI@Meta (2024), Qwen2.5-72B Team (2024), Gemini-1.5-Pro Team et al. (2023), Claude-3.5-Sonnet Anthropic (2024), GPT-4o OpenAI (2024), DeepSeek-v3 Liu et al. (2024). Detailed information of fine-tuning is provided in Appendix F.
@@ 5.3.1 Impact of Hateful Slang on Chinese Hate Speech Detection
LLM APIs demonstrate superior performance on the subset of the test set containing hateful slang. With the exception of a slight decrease in the soft matching metric for argument identification, nearly all other metrics show an improvement. We believe that this is primarily due to two factors: First, hateful slang often appears as the target, and the explicit target terms effectively compensate for the limitations of LLM APIs in span boundary recognition. Second, the identification of hateful slang often requires rich background knowledge; for example, Example 2 in the figure requires knowledge of Chinese history and related common usages, which is precisely where LLM APIs excel.
@@ 6 Conclusions
Our Annotated Lexicon covers a wide range of commonly used Chinese hateful slang; however, due to the dynamic nature and rapid evolution of the Chinese internet language, certain emerging or less widely recognized hateful slang (e.g., more complex word transformations or domain-specific slang) might not be fully captured. Finally, this study focuses primarily on textual features and does not consider non-textual elements, such as images, videos, or metadata about the authors, which may limit the model’s ability to comprehensively detect multimodal hate speech.
In the future, we aim to expand the scope of our dataset to include more diverse contexts and hateful slang, while exploring multimodal and cross-domain hate speech detection methods to improve overall performance and applicability.
@@
To minimize the psychological impact of evaluating harmful content, we have implemented a multi-faceted approach. This includes obtaining informed consent after thoroughly explaining the nature of the content they may encounter, carefully managing exposure by limiting weekly evaluation volumes, and empowering annotators to immediately cease work should they experience any discomfort. Furthermore, we proactively monitor their mental health through regular check-ins, ensuring a supportive and responsible working environment.
@@ Appendix A Details of the Annotated Lexicon
We use JSON files to store the annotated lexicon, which is a mainstream coding specification to facilitate machine-readability. The data structure is defined as Sample = (Term, [Groups], Definition), where Groups indicates the groups that are commonly targeted by the hateful slang. If a term does not have a specific targeted group, then Groups is set to "others".
  • Term: 默 (Mò)
    Meaning: Silence
    Groups: Racism
    Definition: “默”是“黑犬”的拼字,指黑色的狗,对黑人群体的侮辱性称呼。
    Definition in English: The term “默” (Mò, meaning silence) is a spelling variant of “黑犬” (black dog), an insulting term for the Black community.
  • Term: 金针菇
    Meaning: Enoki mushroom (metaphor for male genitalia)
    Groups: Sexism
    Definition: 用于对男性生殖器官的比喻,常含贬义。
    Definition in English: “Enoki mushroom” is a derogatory metaphor for male genitalia.
  • Term: 小仙女
    Meaning: Little fairy (refers to selfish, hypocritical, or annoying women)
    Groups: Sexism
    Definition: 在某些情况下被用来指代那些自私自利、双标、让人厌恶的女性。
    Definition in English: “Little fairy” is used in some contexts to refer to women perceived as selfish, hypocritical, or annoying.
  • Term: 4000+
    Meaning: Homophone for “死全家” (may your whole family die)
    Groups: Others
    Definition: “4000+”是“死全家”中文谐音,常用于恶意辱骂。
    Definition in English: “4000+” is a homophone for “死全家” (may your whole family die), often used in malicious insults.
@@ Appendix E Derivative Rules of Chinese Hateful Slang
Lu et al. (2023) propose seven types of derivative rules: Deformation (separating and combining individual characters), Homophonic (similar pronunciation), Irony (positive words used ironically to insult), Abbreviation (shortening and contracting sensitive words), Metaphor (degrading targets into something sarcastic), Code Mixing (use of non-Chinese codes to emphasize tone), and Borrowed Words (foreign phonetic words with specific toxic cultural connotations). We add two more derivative rules, Historical Allusions and Stereotypes.
@@
Appendix F Detailed Information of the Fine-tuning
We use LLaMA-Factory [^2] https://github.com/hiyouga/LLaMA-Factory for fine-tuning, which is a widely used framework for fine-tuning LLMs. To avoid overfitting, we monitor the trends of training and test losses and confirm through preliminary experiments that the performance stabilizes after 10 epochs. Therefore, all models are trained for 10 epochs. We adopt the LoRA method and evaluate the results.
To reduce hyperparameter sensitivity, we train models for each learning rate and select the result with the highest F1 score on the test set, then calculate the final performance by weighted averaging. The hyperparameters are presented in Table 8.
Hyperparameters Value
Epochs 10
Batch size 2
Learning rate 1e-5, 2e-5, 3e-5,, 4e-5, 5e-5
Cutoff length 1024
Compute type fp16
Gradient accumulation 8
Maximum gradient norm 1.0
Table 8: Annotators Demographics.
Appendix G Detailed Information of the Annotators
To mitigate bias, we assemble a diverse group of annotators, encompassing a variety of genders, ages, ethnicities, regions, and educational backgrounds. The statistical information is presented in Table 9.
Characteristic Demographics
Gender 8 male, 8 female
Age 7 age < 25, 9 age $\geq$ 25
Race 10 Asian, 6 others
Region From 9 different provinces
Education 5 BD, 6 MD, 5 Ph.D.
Table 9: Annotators Demographics.
Appendix H License
We confirm that the dataset is licensed under the Creative Commons Attribution-NonCommercial 4.0 International (CC BY-NC 4.0) license.